OpenStudy (anonymous):

In "slopes", how do i solve y=4/3x+b for b?

OpenStudy (anonymous):

You need another point, like (3, 1) Then, plug in 3 and 1 1 = (4/3)(3) +b and solve for b 1 = 4 + b or b = -3

OpenStudy (anonymous):

I think i get it. I need to plug in one of the coordinates from the line for y and x?

OpenStudy (anonymous):

yes. nice job ... signing out, should be online again in 12-13 hours ... goodnight.

OpenStudy (anonymous):

Goodnight!

OpenStudy (anonymous):

Can someone give me another example?

OpenStudy (anonymous):

...No? Alright...